The grade of a highway up a hill is 20%. How much change in a horizontal distance is there if the vertical height of the hill is 80 feet?

Answer by TimothyLamb(4379)

g = 100 * rise / run
---
g = 100 * rise / run
20 = 100 * 80 / run
20 = 8000 / run
--
run = 8000 / 20
run = 400 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php